OpenStudy (anonymous):

Teri and Jennifer have 480 holiday cookies altogether. Teri has 1 23 as many cookies as Jennifer. How many cookies does Jennifer have?

OpenStudy (anonymous):

how much has tery?

OpenStudy (anonymous):

1 and 2/3 as many cookies as jennifer

OpenStudy (retireed):

Write down what we know first.. Jennifer has J cookies Terry has j times 1.667 cookies (or 3/3 + 2/3 = 5/3) together they have J + J(5/3) = 480 cookies Solve for J. Show me.

OpenStudy (anonymous):

480/5/3?

OpenStudy (anonymous):

=32

OpenStudy (retireed):

No

OpenStudy (anonymous):

but you have to have get j by yourself

OpenStudy (retireed):

check your answer 32 + 32(1.667) is not equal to 480

OpenStudy (anonymous):

oh. i did that completely wrong

OpenStudy (retireed):

32 + 32(1.667) is not equal to 480 equals 85.333 cookies

OpenStudy (retireed):

Yes you did do it completely wrong, but you were right in that you need to get the J by itself. Show me

OpenStudy (anonymous):

do you do the reciprocal of the 5/3?

OpenStudy (retireed):

1 over 5/3. 1 / (5/3) which is 3/5

OpenStudy (anonymous):

I'm so confused

OpenStudy (retireed):

It's an invert kind of thing, like the reciprocal of 2 is 1/2.

OpenStudy (mathmale):

I'd let j=number of cookies belonging to Jenn, and t=number belonging to Teri. Then t = 1.67 j, or t= (5/3) J, or t = (1 2/3)*j. What would the sum of t and j equal, and why?

OpenStudy (anonymous):

I'm still confused

OpenStudy (anonymous):

I'm sorry :(

OpenStudy (retireed):

j + t = 480 J + J (5/3) = 480 Solve for j. How can you isolate the J ? It is confusing at the beginning just hang in there.

OpenStudy (anonymous):

do 3/5 on both sides?

OpenStudy (mathmale):

J=# of coookies Jenn has t=(5/3)j=# of cookies Teri has 480=total number of cookies the girls have If you add t and j together, what does that sum represent? Can you eliminate one of the variables, either t or j?

OpenStudy (retireed):

What?

OpenStudy (mathmale):

RetirEEd wrote the following: j + t = 480 J + J (5/3) = 480 Solve for j. How can you isolate the J ? this is precisely what I was recommending. See if you can find the value of j. Once you have j, find t.

OpenStudy (anonymous):

you take the reciprocal of the 5/3 correct?

OpenStudy (retireed):

No

OpenStudy (anonymous):

do you multiply the 5/3 to 480

OpenStudy (anonymous):

is it 320 for teri

OpenStudy (retireed):

No. What you do to one side of the equation you have to do to the other side. if you multiply 5/3 to 480 you have to multiply it to the J + J (5/3). And that won't isolate the J.

OpenStudy (retireed):

Show me the work first, isolate the J. And tell you if it is right.

OpenStudy (anonymous):

If you want me to get the js by themselves don't I have to do the reciprocal? Cause thats the only way i know how to get the J by itself

OpenStudy (mathmale):

j = # cookies Jennifer owns t= # cookies teri owns = (5/3)j j+t=j + (5/3)j =480 cookies, total. Solve for j. then find t.

OpenStudy (retireed):

mathmale technically they only asked for the number of cookies Jennifer has.

OpenStudy (anonymous):

I still don't understand what I'm supposed to do if i can't do the reciprocal

OpenStudy (retireed):

but it is good to solve for both and add them together to check your answers

OpenStudy (retireed):

Last clue....... The left hand equation has a J in both terms. Factor out the Js and divide both sides by the what is left in the parenthesis. Have a good night. Time to watch my Cowboys loose.